OpenStudy (anonymous):

find 3 consecutive odd intergers whose sum is 249

OpenStudy (anikhalder):

Let the first odd no be (a-2), 2nd consecutive odd no is then (a-2)+2=a and the third consecutive add no is (a+2). so the 3 consecutive odd nos are : a-2, a ,a+2 For example, 3 is an odd no. (which is like a) The consecutive odd nos. before and after 3 are: 3-2=1 which is like (a-2) and 3+2=5 and this one is like (a+2) Hence,proceeding with the question which says the summation of the 3 consecutive odd nos give 249, (a-2)+a+(a+2)=249 Simplifying, a-2+a+a+2=249 =>3a=249 =>a=83 Therefore, a=83 a-2=81 a+2=85 Hence the 3 consecutive odd nos are: 81,83,85